OpenStudy (anonymous):

Find the slope of the line passes through (-8,-3) and (2,3). Find the slope of the line passes through (-8,-3) and (2,3). @Mathematics

OpenStudy (anonymous):

Slope: \[ \frac{2+8}{3+3}=\frac{5}{3} \]

OpenStudy (anonymous):

Actually, it's 3/5, but I think you knew that.